# Supporting Bipolar Fabry–Pérot charge interferometer in periodically electron-irradiated graphene

### v1

@@ Abstract
Electron optics deals with the wave-nature of charge carriers to induce, investigate and exploit coherent phenomena in solid state devices, in analogy with optics and photonics. Typically, these goals are achieved in complex electronic devices taking advantage of the macroscopically coherent charge transport in two dimensional electron gases and superconductors. Here, we demonstrate a simple counterintuitive architecture employing intentionally-created lattice defects to induce collective coherent effects in the charge transport of graphene. More specifically, multiple Fabry–Pérot cavities are produced by irradiating graphene via low-energy electron-beam to form periodically alternated defective and pristine nano-stripes. The enhanced hole-doping in the defective stripes creates potential barriers behaving as partially reflecting mirrors and resonantly confining the carrier-waves within the pristine areas. The interference effects are both theoretically and experimentally investigated and manifest as sheet resistance oscillations up to 30 K for both polarities of charge carriers, contrarily to traditional electrostatically-created Fabry-Pérot interferometers. Our findings propose defective graphene as an original platform for the realization of innovative coherent electronic devices with applications in nano and quantum technologies.
Keywords: Electron-optics, quantum interference, defective graphene
@@ 1 Introduction
Electron optics exploits the wave nature of the electrons in conductors to explore fundamental properties of charges in crystals. Innovative solid state devices and electronic circuits can be developed, in analogy to optical devices and quantum optics experiments, for applications in single-electron quantum optics, nanoelectronics and quantum technologies [1, 2, 3]. Inspired by the optical counterpart, electron optical elements, such as beam splitters [1, 4], focusing lenses [5, 6] and waveguides [1, 7], have been developed to confine and manoeuvre the electron waves. Quantum interference is one of the prototypical manifestations of the wave-like nature of electrons. This phenomenon occurs when two or more electron waves superimpose while coherently travelling in the same conductor. Usually, interference effects are generated in devices where the electrons are spatially confined along specific directions, thus enabling the formation of optics-like interferometers having, for instance, Fabry-Pérot [8, 9, 10] or Mach-Zehnder [11, 12] geometry. Indeed, traditional electronic interferometers have been realized in low dimensional systems [11, 13, 14, 15, 8] or superconductors [16, 17].
The advent of graphene has opened new paths towards novel electron wave devices [18, 19, 20, 21]. Compared to conventional two-dimensional electron gases, graphene offers easily tunable electrical properties, coherent transport accessible up to high temperatures [22], as well as negative refraction of charge carriers [23]. Several optics-like electronic elements, such as interferometers [24, 25], beam splitters [4] and electron-lenses [26, 27], have been currently realized. The key element in graphene coherent devices is the external control of the local electrochemical potential, that allows the formation of energy barriers at the interface of regions of different doping (junction). In the analogy with optics, the electrochemical potential plays the same role of the refractive index in ray optics when considering the semiclassical description [23], whereas the energy barriers act as semitransparent mirrors. Indeed, graphene potential barriers have electron transmission probability () that depends on the junction sharpness and the incident angle (Snell’s law) [28, 29]. In particular, graphene shows Klein tunneling, that is = 1 for orthogonal incidence [20, 29]. Based on these properties, Fabry-Pérot-like resonant effects have been predicted [29] and observed in single or multiple gate graphene field-effect transistors. In particular, these phenomena manifest in the form of gate-induced oscillations in the total charge current flowing along the structure [25, 19, 30, 31, 32, 33, 24, 34]. Indeed, when sweeping the gate voltage (i.e., the electrochemical potential), the Fermi wavelength of the charge carriers is tuned. When the Fermi wavelength matches the wavelengths of the allowed cavity modes, the charge carriers are spatially confined, unlocking resonant phenomena.
@@
Here, we demonstrate a simple charge carrier interferometer architecture based on the succession of pristine and defective graphene nano-stripes (Figure 1a). In particular, our approach makes use of artificially-created defects in the graphene lattice via electron-beam irradiation to generate charge carrier quantum interference with clear fingerprints in the device charge carrier transport. In other words, typically-undesired defects are here exploited to induce and regulate coherent phenomena in the charge carrier transport of graphene sheets, thus refuting the parallelism between defects and disorder. Moreover, our approach is based on direct writing. It exploits the standard electron-beam lithography system, but it does not involve polymeric resist materials, thus reaching a cavity length scale that is not achievable via standard electrostatic-gating-based approaches.
In our structure, the defects form the mirrors of a FP charge carrier cavity by introducing controlled potential barriers (Figure 1b). Indeed, air exposed defective graphene can be strongly hole doped and operate as potential barrier. Our FP interferometer is realized by regularly alternating parallel defective and quasi-ballistic nano-stripes with nanometric lateral resolution [8]. When tuning the electrochemical potential, periodic oscillations in the sheet resistance of graphene are observed both for holes and electrons, demonstrating the bipolar operation regime. The FP modes are affected by possible disorder in the potential barrier heights, as confirmed by our simulations based on the Landauer-Büttiker formalism. Indeed, the simulations reproduce both the oscillatory behavior of the sheet resistance and the formation of multi-peak resonant modes in the presence of sizeable barrier inhomogeneity. Finally, temperature-dependent measurements show that the FP modes are robust up to 30 K, thus allowing us to extract the charge carrier coherence time, fundamental for the interferometer operation.
@@ 3 Conclusions
We presented an original approach to generate coherent quantum phenomena in graphene sheets by exploiting the wave nature of charge carriers. In particular, electron-irradiation-induced defects in the graphene lattice are exploited to mediate the coherent carrier transport by acting as partially transmitting mirrors, that confine the charge carriers within pristine quasi-ballistic nano-stripes (resonant structures). The regular alternation of defective and pristine nano-stripes creates independent FP cavities, whose interference fringes manifest as thermally-robust (up to 30 K) oscillations in the graphene sheet resistance. These oscillations are observed in series of FP interferometers having different cavity lengths. Moreover, the charge carrier interference patterns are present for both electrons and holes, revealing a bipolar behaviour, that is typically hardly observable in gate-induced barriers. As confirmed by numerical simulations, possible disorder in the barrier height generate multi-peak main FP modes. Finally, the interference pattern is not correlated to the number of the potential barriers, indicating that the FP cavities are independent.
Differently from the standard electrostatic approach based on gating, this methodology is extremely simple. On the one hand, it avoids the use of metal stripes that limit the minimum cavity length (typically 50 nm), offering the possibility to fully scale down and produce resonant structures also on a chip scale-level (if using high quality large-area graphene). On the other hand, thanks to the enhanced chemical reactivity, it offers an additional knob to control the resonant cavities: when properly functionalized [40], the defective graphene allows to define ad hoc the properties of the potential barriers. Consequently, electron-irradiated graphene represents an original route to develop innovative coherent electronics devices for the manipulation of carrier waves and the manifestation of quantum charge-optics effects.
@@ 4 Experimental Section
Supporting Information
Supporting Information is available from the Wiley Online Library or from the author.
@@ Acknowledgements
Supporting information:
Bipolar Fabry–Pérot charge interferometer in periodically electron-irradiated graphene

### v2

# Tailoring coherent charge transport in graphene by deterministic defect generation

@@ Abstract
Harnessing the wave-nature of charge carriers in solid state devices, electron optics investigates and exploits coherent phenomena, in analogy with optics and photonics. Typically, this requires complex electronic devices leveraging macroscopically coherent charge transport in two-dimensional electron gases and superconductors. Here, collective coherent effects are induced in a simple counterintuitive architecture by defect engineering. Deterministically introduced lattice defects in graphene enable the phase coherent charge transport by playing the role of potential barriers, instead of scattering centres as conventionally considered. Thus, graphene preserves its quasi-ballistic quantum transport and can support phase-matched charge carrier-waves. Based on this approach, multiple electronic Fabry–Pérot cavities are formed by creating periodically alternating defective and pristine nano-stripes through low-energy electron-beam irradiation. Indeed, defective stripes behave as partially reflecting mirrors and resonantly confine the charge carrier-waves within the pristine areas, giving rise to Fabry–Pérot resonant modes. These modes experimentally manifest as sheet resistance oscillations, as also confirmed by Landauer-Buttiker simulations. Moreover, these coherent phenomena survive up to 30 K for both polarities of charge carriers, contrarily to traditional monopolar electrostatically created Fabry-Pérot interferometers. Our study positions defective graphene as an innovative platform for coherent electronic devices, with potential applications in nano and quantum technologies.
Keywords: Defective graphene, coherent charge transport, quantum interference.
@@ 1 Introduction
Coherent charge transport exploits the wave nature of the electrons in conductors to explore fundamental properties of charges in crystals. Innovative solid state devices and electronic circuits can be developed, in analogy to optical devices and quantum optics experiments, for applications in single-electron quantum optics, nanoelectronics and quantum technologies [1, 2, 3, 4, 5, 6, 7]. Quantum interference is one of the prototypical manifestations of the wave-like nature of electrons. Usually, interference effects are generated in devices where the electrons are spatially confined along specific directions, thus enabling the formation of optics-like interferometers having, for instance, Fabry-Pérot [8, 9, 10] or Mach-Zehnder [11, 12] geometry. Indeed, traditional electronic interferometers have been realized in low dimensional systems [11, 13, 14, 15, 8] or superconductors [16, 17].
The advent of graphene has opened new paths towards novel electron wave devices [18, 19, 20, 21]. Compared to conventional two-dimensional electron gases, graphene offers easily tunable electrical properties and coherent transport accessible up to high temperatures [22]. Several optics-like electronic elements have been currently realized [4, 23, 24, 25, 26]. The key element in graphene coherent devices is the external control of the local electrochemical potential, that allows the formation of energy barriers at the interface of regions of different doping (junction). In the analogy with optics, the electrochemical potential plays the same role of the refractive index in ray optics when considering the semiclassical description [27], whereas the energy barriers act as semitransparent mirrors. Indeed, graphene potential barriers have electron transmission probability () that depends on the junction sharpness and the incident angle (Snell’s law) [28, 29]. In particular, graphene shows Klein tunneling, that is = 1 for orthogonal incidence [20, 29]. Based on these properties, Fabry-Pérot-like resonant effects have been predicted [29] and observed in single or multiple gate graphene field-effect transistors. In particular, these phenomena manifest in the form of gate-induced oscillations in the total charge current flowing along the structure [24, 19, 30, 31, 32, 33, 23, 34]. Indeed, when sweeping the gate voltage (i.e., the electrochemical potential), the Fermi wavelength of the charge carriers is tuned. When the Fermi wavelength matches the wavelengths of the allowed cavity modes, the charge carriers are spatially confined, unlocking resonant phenomena.
@@
Here, we demonstrate collective coherent effects induced in graphene by defect engineering. To this scope, we produced a simple charge carrier interferometer architecture based on the succession of pristine and defective graphene nano-stripes (Figure 1a). Our approach makes use of artificially-created defects in the graphene lattice via electron-beam irradiation to generate charge carrier quantum interference with clear fingerprints in the device charge carrier transport. In other words, typically-undesired defects are here exploited to induce and regulate coherent phenomena in the charge carrier transport of graphene sheets, thus refuting the parallelism between defects and disorder. Moreover, our approach is based on direct writing. It exploits the standard electron-beam lithography system, but it does not involve polymeric resist materials, thus reaching a cavity length scale that is not achievable via standard electrostatic-gating-based approaches. Contrarily to the typical elastic scattering-induced localization effects obtained in disordered conductors [38], in our structure graphene defects simply introduce controlled potential barriers (Figure 1b), while preserving the quasi-ballistic transport in the defect-free areas. Indeed, air exposed defective graphene can be strongly hole doped, thus operating as potential barrier. Consequently, our structure implements a FP charge carrier cavity by regularly alternating parallel defective, acting as mirrors, and quasi-ballistic nano-stripes with nanometric lateral resolution [7]. When tuning the electrochemical potential, periodic oscillations in the sheet resistance of graphene are observed both for holes and electrons, demonstrating the bipolar operation regime. The FP modes are affected by possible disorder in the potential barrier heights, as confirmed by our simulations based on the Landauer-Büttiker formalism. Indeed, the simulations reproduce both the oscillatory behavior of the sheet resistance and the formation of multi-peak resonant modes in the presence of sizeable barrier inhomogeneity. Finally, temperature-dependent measurements show that the FP modes are robust up to 30 K, thus allowing us to extract the charge carrier coherence time, fundamental for the interferometer operation.
@@ 3 Conclusions
We presented an original approach to generate coherent quantum phenomena in graphene sheets by exploiting the wave nature of charge carriers. In particular, electron-irradiation-induced defects in the graphene lattice are exploited to mediate the coherent carrier transport by acting as partially transmitting mirrors, that confine the charge carriers within pristine quasi-ballistic nano-stripes (resonant structures). This capability strongly differs from the conventional role of lattice defects as elastic scattering centres in diffusive conductors, where localization phenomena usually occur and dominate the charge transport. The regular alternation of defective and pristine nano-stripes creates independent FP cavities, whose interference fringes manifest as temperature-robust (up to 30 K) oscillations of the graphene sheet resistance. Moreover, these charge carrier interference patterns are present for both electrons and holes, revealing a bipolar behaviour, which is typically hardly observable in gate-induced barriers. As confirmed by numerical simulations, possible disorder in the barrier height generates multi-peak FP modes resulting in an increased number of resistance oscillations. Finally, the interference pattern is not correlated to , indicating that thus the FP cavities are independent. Differently from the standard electrostatic approach based on gating, this methodology is extremely simple. On the one hand, it avoids the use of metal stripes that limit the minimum cavity length (typically 50 nm), offering the possibility to fully scale down and produce resonant structures also on a chip scale-level (if using high quality large-area graphene). On the other hand, thanks to the enhanced chemical reactivity, it offers an additional knob to control the coherent charge transport: when properly functionalized [40], the defective graphene allows to define ad hoc the properties of the potential barriers. Consequently, electron-irradiated graphene represents an original route to develop innovative coherent electronics devices for the manipulation of carrier waves and the manifestation of quantum charge-optics effects.
@@ Acknowledgements
This material is based in part upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Superconducting Quantum Materials and Systems Center (SQMS) under contract number DE-AC02-07CH11359.
@@
Data availability
The data that support the findings of this study are available in this article and its Supporting Information. Any other relevant data are available from the corresponding authors upon reasonable request.
Contributions
N.M. fabricated the devices and carried out the micro-Raman characterization. N.M. and F.P. performed charge carrier transport experiments with inputs from F.B.. P.M. and M.M. conducted numerical simulations. N.M., F.P., S.R., A.T., and F.B. analysed and interpreted the data. F.B. conceived the idea of the project, coordinated research efforts and wrote the paper. All authors discussed the data and participated in preparing the manuscript.
SUPPORTING INFORMATION